What is the area of a rectangle whose length is four more than four times its width if the perimeter is 42
Answer by ReadingBoosters(3246)

a = lw
p = 2l+2w = 42
l=4w+4
Substitute
2(4w+4) + 2w = 42
8w + 8 + 2w = 42
10w = 48 - 8
10w = 40
width = 4
length = 4(4)+4 = 20
...
a = lw
a = 4(20) = 80
**be sure to add units**
..................
Delighted to help.
-Reading Boosters
Wanting for others what we want for ourselves.